# All-in-one Graph-based Indexing for Hybrid Search on GPUs

### v1

@@ Abstract.
Hybrid search has emerged as a promising paradigm to overcome the limitations of single-path retrieval, enhancing accuracy for applications like recommendations, information retrieval, and Retrieval-Augmented Generation. However, existing methods are constrained by a trilemma: they sacrifice flexibility for efficiency, suffer from accuracy degradation due to separate retrievals, or incur prohibitive storage overhead for flexible combinations of retrieval paths.
@@
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/xxx.
@@ 1. Introduction
Recent advancements in vector databases have substantially improved the accuracy and efficiency of dense vector retrieval (Pan et al., 2024; Gao and Long, 2024; Cai et al., 2024). State-of-the-art approximate nearest neighbor search algorithms now consistently achieve over 99% recall for top-$k$ neighbor searches. However, the end-to-end accuracy—that is, the accuracy of retrieved documents rather than vector similarity—remains limited. This is because the minimal distance between query and answer vectors in the embedding space does not guarantee their semantic relevance in natural language (Zhang et al., 2024; Sawarkar et al., 2024; Liu et al., 2025). This fundamental discrepancy between geometric proximity and semantic meaning hinders the broader adoption of vector databases in critical areas such as search engines (Li et al., 2014, 2018; Gou et al., 2024), recommendation systems (Miao et al., 2022; Parchas et al., 2020; Yang et al., 2024), and Retrieval-Augmented Generation (RAG) (Ang et al., 2024; Yu et al., 2025; Chen et al., 2025a; Jiang et al., 2024).
@@
In addition to the popular use of dense vectors, alternative retrieval methods utilize statistics-based (Wang et al., 2017a; Mallia et al., 2017) and learned-based (Kong et al., 2023; Formal et al., 2024b; Bruch et al., 2024b) sparse vectors to improve the end-to-end accuracy. While these sparse vectors offer superior interpretability and cross-domain robustness, their semantic representation capabilities are generally weaker than dense vectors. Consequently, dependence on any single retrieval path is often insufficient for achieving high end-to-end retrieval relevance (Weller et al., 2025; Mala et al., 2025). To overcome this limitation, hybrid search has emerged as a promising solution (Ouyang et al., 2025; Mala et al., 2025; Wang et al., 2025). As illustrated in Figure 1, two primary methodologies have been proposed: (1) Fusion retrieval integrates dense and sparse vectors by mapping them into a unified space via dimensionality reduction (Bruch et al., 2024a) or by constructing the graph-based index with weighted distance calculations (Zhang et al., 2024). However, these approaches often suffer from low efficiency and precision, or exhibit poor extensibility. Furthermore, they are typically limited to dense and sparse vectors, failing to accommodate the complex requirements of real-world applications. (2) Multi-path retrieval constructs separate indexes for various retrieval paths, including dense vector (Malkov and Yashunin, 2018; Fu et al., 2019; Yang et al., 2025b), sparse vector (Kong et al., 2023; Formal et al., 2024b; Bruch et al., 2024b), and full-text search (Wang et al., 2017a; Mallia et al., 2017). For a given query, this paradigm retrieves the top-$k^{\prime}$ neighbors from each path independently. These intermediate results are subsequently fused using re-ranking methods—such as Reciprocal Rank Fusion (RRF) (Cormack et al., 2009; Bruch et al., 2023), Weighted Sum (inf, 2025; mil, 2025b), or ColBERT (Formal et al., 2024a; Khattab and Zaharia, 2020)—to produce the final top-$k$ ($k\leq k^{\prime}$) list. Due to its effectiveness and flexibility, this paradigm is widely adopted in modern databases (mil, 2025a; inf, 2025; wea, 2025; ela, 2025).
@@
Despite the effectiveness, the paradigm of multi-path retrieval introduces two primary problems due to its decoupled architecture. (1) Index storage overhead. Each retrieval path necessitates the construction, maintenance, and storage of a dedicated index. This not only increases system complexity but also incurs significant storage and operational overhead [^1] For instance, Infinity (inf, 2025) requires 5GB to store the indexes of three retrieval paths for a dataset with 1M documents, where the HNSW index size for dense vector retrieval only accounts for 23%. (Zhang et al., 2024; Liang et al., 2025). (2) Retrieval efficiency and accuracy. The optimal results for a hybrid query may not be present within the top-$k^{\prime}$ results from any single path (Wang et al., 2021b, 2024a; Zhu et al., 2024), leading to decreased accuracy. As a result, neither Fusion retrieval nor Multi-path retrieval can simultaneously deliver high accuracy, efficiency, and extensibility.
To overcome the aforementioned limitations, inspired by Edgar Allan Poe’s literary theory of ”Unity of Effect”, this work explores unifying diverse multi-path indexes into a single, all-in-one hybrid index, leveraging the GPU to achieve efficient construction and real-time retrieval. The graph-based index has recently gained prominence as a leading approach for approximate nearest neighbor search (ANNS), owing to its superior efficiency and accuracy (Voruganti and Özsu, 2025; Jiang et al., 2025; Xu et al., 2024; Wang et al., 2021a). This naturally raises a key research question: how can we design a unified graph-based index for effective hybrid search? However, such a design is non-trivial. Practical applications with real-time requirements—such as recommendation systems, search engines, and RAG—demand a hybrid index that is simultaneously efficient, effective, and flexible. We identify two primary challenges that must be addressed:
@@
Challenge I: How to design a unified and flexible graph-based hybrid index? The fundamental heterogeneity of retrieval methods presents the primary obstacle: dense vectors utilize graph-based indexes, sparse vectors and full-text search rely on inverted indexes, and knowledge graphs employ entity-relationship structures. Fusing these disparate architectures into a single, coherent index without compromising performance is non-trivial. Furthermore, optimal fusion weights for the similarity from each retrieval path are inherently dynamic, varying by query context and user preference. Pre-computing and storing indexes for all possible weight combinations is infeasible. Finally, exhaustive use of all available paths for every query is suboptimal, unnecessarily compromising efficiency when fewer paths would suffice. Consequently, designing a flexible structure that supports arbitrary path combinations without index reconstruction remains a critical challenge.
@@
Challenge II: How to achieve efficient construction and effective search on the hybrid index? While Inner Product (IP) serves as the primary similarity metric (Bruch et al., 2024a; Wang et al., 2025), existing graph indexes optimized for it suffer from inefficient construction and redundant edge connectivity. The indexing and querying processes also involve massive, hybrid distance calculations. For example, dense vectors and sparse vectors comparisons require vector dot product and set intersection operations, respectively. These disparate computational characteristics create challenges for efficient memory access and effective GPU parallelization. Furthermore, the computational demands of edges from various retrieval paths in the hybrid index introduce multiple efficiency bottlenecks. Semantic matching alone proves inadequate for complex reasoning tasks. Although knowledge graphs provide complementary logical similarity (Zhou et al., 2025; Cao et al., 2025), a fundamental granularity mismatch exists: our index operates on document-level representations while knowledge graphs model fine-grained entity relationships. This disparity complicates the effective integration of logical reasoning into the vector search process.
@@
To address these challenges, we propose Allan-Poe, a unified and flexible graph-based hybrid index accelerated by GPUs. Our solution integrates dense, sparse, full-text, and knowledge graph-based retrieval through an isolated heterogeneous edge storage mechanism. This design supports any combination of retrieval paths without requiring index reconstruction or sacrificing efficiency. We create a Unified Semantic Metric Space that fuses multiple vector representations from diverse retrieval paths into a single similarity metric, with theoretical proof demonstrating its effectiveness for arbitrary fusion weights. To achieve efficient hybrid index construction, we develop a GPU-accelerated indexing pipeline featuring: (1) a warp-level hybrid distance computation kernel optimizing both dense and sparse operations parallelized on GPUs; (2) RNG-IP joint pruning that maintains search quality while reducing index complexity by combining Relative Neighborhood Graph (RNG) and Inner Product (IP) neighbor pruning; (3) keyword-aware neighbor recycling that preserves keyword search functionality by efficiently recycling the pruned neighbors to ensure keyword-based navigation on the index; and (4) logical edge augmentation that integrates the entity-level knowledge graph edges into the document-level hybrid index. To deliver a high-accuracy and high-throughput search service, we design a dynamic query framework on GPUs, incorporating: (1) dynamic heterogeneous edges loading for efficient traversal on the hybrid index; and (2) entity-document joint traversal for knowledge graph integration.
@@
  • We analyze the limitations of existing retrieval paradigms and derive a set of design principles for effective and flexible hybrid indexes (§ 2).
  • We integrate multi-path retrieval into a unified semantic metric space, demonstrating its capability to handle fused distances with arbitrary weights. Based on this foundation, we design a hybrid index that supports any combination of retrieval paths without requiring reconstruction (§ 3).
  • We propose an efficient GPU-accelerated framework for index construction and query processing. This framework enhances the inner product neighbors search, optimizes the hybrid distance calculation, and seamlessly integrates semantic and logical similarities (§ 4).
  • We conduct comprehensive experiments on 6 real-world datasets, demonstrating that Allan-Poe achieves superior performance compared to state-of-the-art methods (§ 5).
@@ 2.1. Related Work
(4) Knowledge graph search implements logical and semantic retrieval by converting queries into subgraphs and identifying similar structures through subgraph matching algorithms (Hu et al., 2017; Yuan et al., 2021; Sun et al., 2022). Subsequent enhancements incorporate entity and relation embeddings for improved efficiency (Zheng et al., 2020; Wang et al., 2017b). Recent approaches like GraphRAG leverage dense vector search and community detection for document retrieval of global questions (Edge et al., 2024; Cao et al., 2025; Zhou et al., 2025; Li et al., 2025a). While GraphRAG excels at summarization tasks, it typically underperforms vanilla RAG for question answering (Han et al., 2025). Unlike GraphRAG, our work selectively integrates logical information from knowledge graphs to enhance vector search, establishing a distinct paradigm applicable to more general scenarios beyond RAG.
@@
(1) Fusion retrieval integrates multiple retrieval methods within a unified index structure. Current fusion methods are primarily restricted to two-path combinations (Wang et al., 2025). For instance, DS-ANN (Zhang et al., 2024) employs pre-defined fusion weight to combine dense and sparse vectors, constructing an HNSW index (Malkov and Yashunin, 2018) for efficient querying. While efficient, this method requires complete index reconstruction if the fusion weights change. IVF-Fusion (Bruch et al., 2024a) addresses this by reducing the dimensionality of sparse vectors before combining them with dense vectors, then using an IVF index for retrieval. Although this eliminates weight-dependent reconstruction, it limits the flexibility to select different retrieval paths for varying scenarios.
(2) Multi-path retrieval represents a more flexible paradigm that executes searches separately across different indexes and subsequently fuses the results (mil, 2025a; inf, 2025; ela, 2025). However, as discussed in Section 1, this flexibility adversely affects query efficiency and accuracy while complicating index management.
@@ 2.2. Motivation
Recently, dense vector search has become the most popular paradigm in vector databases among single-path retrieval methods, distinguished by its capacity for comprehensive semantic representation (Pan et al., 2024; Wang et al., 2021a). Numerous specialized indexes have been developed to enhance their query efficiency and accuracy (Gao et al., 2025; Gou et al., 2025; Yin et al., 2025; Peng et al., 2023). However, vector similarity alone does not guarantee end-to-end relevance between queries and documents. To investigate this limitation empirically, we conduct experiments using two established real-world QA datasets: NaturalQuestions (NQ) (Wang et al., 2024b) for simple question answering and 2WikiMultiHopQA (WM) (Ho et al., 2020) for multi-hop question answering. For each dataset, we evaluate the first 1,000 queries, each associated with 1-3 ground-truth documents. We employ the BGE-M3 model (Chen et al., 2024) for the embedding of both documents and queries. The ground-truth baselines of vector similarity are established by computing the top-10 nearest neighbors via brute-force vector similarity search. We then assess the retrieval accuracy of two state-of-the-art graph-based indexes: HNSW (CPU-based) (Malkov and Yashunin, 2018) and CAGRA (GPU-based) (Ootomo et al., 2024).
@@
As shown in Figure 2, while vector-similarity accuracy can easily reach 99% within 4ms on the CPU and 0.1ms on the GPU, the corresponding end-to-end accuracy is substantially lower. In practical document retrieval systems, it is this end-to-end accuracy—not vector-similarity accuracy—that dictates performance in downstream tasks. Furthermore, Figure 2 reveals that the end-to-end accuracy for the WM dataset is lower than that for NQ, whereas their vector-based accuracies are comparable. This discrepancy underscores the limitations of dense vector search in handling complex queries. Consequently, reliance on any single-path retrieval method is insufficient for achieving satisfactory end-to-end performance, thereby restricting its utility in downstream applications.
@@
Hybrid search has emerged as a powerful strategy to mitigate the limitations of single-path retrieval and improve end-to-end accuracy (inf, 2025; Wang et al., 2025). To evaluate its effectiveness, we measure retrieval quality using Infinity (inf, 2025), a modern database featuring efficient hybrid search. To better assess the quality of retrieved documents, we employ Normalized Discounted Cumulative Gain at rank $k$ (nDCG@kk(Wang et al., 2013) with $k=5$, which evaluates both the recall and positional ranking of relevant documents in the retrieved results.
@@
Figure 3. Comparison of various retrieval paths using Infinity (inf, 2025). DVS, SVS, and FTS denote dense vector, sparse vector, and full-text search, respectively.
@@
However, no single configuration is optimal for all scenarios. Retrieval with more paths does not consistently outperform fewer or single paths. For example, in Figure 3(a), the two-path combination of dense vector and full-text search yields lower nDCG than the single-path methods using either dense or sparse vectors alone on dataset NQ. And the three-path combination shows lower nDCG than the two-path retrieval of dense and sparse vector search. On other datasets, the accuracy ranking may be completely different. Furthermore, different path combinations present a distinct trade-off between accuracy and efficiency. Although the single-path methods typically achieve lower accuracy than the multi-path approaches, they are more efficient and incur less overhead (Figure 3(b)), sometimes making it preferable for a few real-time applications. For instance, the sparse vector retrieval method achieves comparable accuracy with the three-path combination while maintaining lower latency. These findings underscore the necessity for flexible path combination in hybrid search systems.
@@
Multi-path retrieval is widely adopted for flexible hybrid search in modern vector databases (inf, 2025; mil, 2025a; ela, 2025). Although it achieves superior accuracy compared to single-path approaches, this comes at the cost of increased time overhead. Moreover, this paradigm performs separate retrievals across individual indexes before fusing the results. The optimal results for a hybrid query may not be present within the top-$k$ results from any single path or may be excluded after fusing the results (Wang et al., 2024a, 2021b). To illustrate this, we examine the retrieval of top-5 documents from the NQ dataset using dense and sparse vectors independently. Figure 4 presents a representative example. Using dense vectors, ground-truth documents doc2 and doc4 are ranked 2nd and 30th, respectively, while sparse vectors rank them 3rd and 12th. If we fuse the top-5 results from both paths using equal weights (i.e., $0.5\timesdense$ $0.5\timessparse similarity$), doc4 will be excluded because it only involves the similarity score from the sparse path. To include doc4, more candidates should be included for each retrieval path (e.g., top-30). However, to ensure the accuracy of more candidates, more time is required for each retrieval path, leading to high end-to-end retrieval overhead.

### v2

@@ Abstract.
Hybrid search has emerged as a promising paradigm that combines lexical and semantic retrieval, enhancing accuracy for applications such as recommendations, information retrieval, and Retrieval-Augmented Generation. However, existing methods are constrained by a trilemma: they sacrifice flexibility for efficiency, suffer from accuracy degradation, or incur prohibitive storage overhead for flexible combinations of retrieval paths.
@@
The source code, data, and/or other artifacts have been made available at https://github.com/ZJU-DAILY/Allan-Poe.
@@ 1. Introduction
Recent advancements in vector databases have substantially improved the accuracy and efficiency of dense vector retrieval (Pan et al., 2024; Gao and Long, 2024; Cai et al., 2024). State-of-the-art approximate nearest neighbor search algorithms now consistently achieve over 99% recall for top-$k$ neighbor searches. However, the end-to-end accuracy—that is, the accuracy of retrieved documents rather than vector similarity—remains limited, lagging behind the vector recall by 10%-30% (§2.2). This discrepancy arises because the minimal distance between query and answer vectors in the embedding space does not guarantee their semantic relevance in natural language (Zhang et al., 2024; Sawarkar et al., 2024; Liu et al., 2025). While dense vector retrieval excels at capturing broad semantic similarity, its inherent loss of fine-grained lexical details restricts the accuracy, hindering the broader adoption of vector databases in critical areas such as search engines (Li et al., 2014, 2018; Gou et al., 2024), recommendation systems (Miao et al., 2022; Parchas et al., 2020; Yang et al., 2024), and Retrieval-Augmented Generation (RAG) (Ang et al., 2024; Yu et al., 2025; Chen et al., 2025a; Jiang et al., 2024).
@@
In addition to the popular use of dense vectors, alternative retrieval methods utilize lexicality-based (Wang et al., 2017a; Mallia et al., 2017) and learned-based (Kong et al., 2023; Formal et al., 2024b; Bruch et al., 2024b) sparse vectors to improve the semantic relevance. While these sparse vectors offer superior interpretability and cross-domain robustness, their semantic representation capabilities are generally weaker than dense vectors. Consequently, dependence on any single retrieval path is often insufficient for achieving high end-to-end accuracy (Weller et al., 2025; Mala et al., 2025). To overcome this limitation, hybrid search has recently emerged as a promising solution (Ouyang et al., 2025; Mala et al., 2025; Wang et al., 2025). As illustrated in Figure 1 (a) and (b), two primary methodologies have been proposed: (1) Fusion retrieval integrates dense and sparse vectors by mapping them into a unified space via dimensionality reduction (Bruch et al., 2024a) or by constructing the graph-based index with weighted distance calculations (Zhang et al., 2024). However, these approaches often suffer from low efficiency or exhibit poor flexibility for arbitrary fusion weights. Furthermore, they are limited to dense and sparse vectors, failing to accommodate the complex requirements of real-world applications. (2) Multi-path retrieval constructs separate indexes for various retrieval paths (Wang et al., 2025), including dense vector (Malkov and Yashunin, 2018; Fu et al., 2019; Yang et al., 2025b), sparse vector (Kong et al., 2023; Formal et al., 2024b; Bruch et al., 2024b), and full-text search (Wang et al., 2017a; Mallia et al., 2017). For a given query, this paradigm retrieves the top-$k^{\prime}$ neighbors from each index independently. These intermediate results are subsequently fused using re-ranking methods, such as Reciprocal Rank Fusion (RRF) (Cormack et al., 2009; Bruch et al., 2023), Weighted Sum (inf, 2025; mil, 2025b), or ColBERT (Formal et al., 2024a; Khattab and Zaharia, 2020), to produce the final top-$k$ ($k\leq k^{\prime}$) list. Due to its effectiveness and flexibility, this paradigm is widely adopted in modern databases (mil, 2025a; inf, 2025; wea, 2025; ela, 2025).
@@
Despite the effectiveness of the multi-path retrieval paradigm, it introduces two issues due to the decoupled architecture. (1) Index storage overhead. Each retrieval path necessitates the construction, storage, and maintenance of a dedicated index. This not only increases system complexity but also incurs significant storage overhead [^1] For instance, Infinity (inf, 2025) requires 5GB to store the indexes of three retrieval paths for a dataset with 1M documents, where the HNSW index size for dense vector retrieval only accounts for 23%. (Zhang et al., 2024; Liang et al., 2025). (2) Retrieval accuracy. The optimal results for a hybrid query may not be present within the top-$k^{\prime}$ results from any single path (Wang et al., 2021b, 2024a; Zhu et al., 2024), leading to decreased accuracy.
Collectively, these limitations constrain existing paradigms within a trilemma of flexibility, efficiency, and accuracy. In this paper, we argue that resolving this trilemma requires rethinking hybrid retrieval from a structural perspective. We observe that heterogeneous retrieval paths can be naturally modeled as complementary relationships over a unified space with arbitrary weights, which motivates us to unify diverse retrieval paths within an all-in-one hybrid index. As illustrated in Figure 1(c), this unified architecture significantly reduces storage overhead. Furthermore, it enhances query performance by integrating three mainstream retrieval paths into a single index while optionally leveraging logical structures from knowledge graphs to augment retrieval for complex queries. Constructing such an integrated framework requires careful design. Given the superior efficiency and accuracy of the graph-based index in vector search (Voruganti and Özsu, 2025; Jiang et al., 2025; Xu et al., 2024; Wang et al., 2021a), it represents a promising foundation for building the unified framework. However, incorporating these heterogeneous retrieval paths into a graph index is non-trivial. We identify two primary challenges that must be addressed:
@@
Challenge I: How to design a flexible graph-based hybrid index for hybrid search? The fundamental heterogeneity of retrieval methods presents the primary obstacle: dense vectors utilize graph-based indexes, sparse vectors and full-text search rely on inverted indexes, and knowledge graphs employ entity-relationship structures. Fusing these disparate architectures into a single graph-based index without compromising performance is non-trivial. Furthermore, optimal fusion weights for the similarity from each retrieval path are inherently dynamic, varying by query context and user preference. Pre-computing and storing indexes for all possible weight combinations is infeasible. Finally, exhaustive use of all available paths for every query is suboptimal, which unnecessarily diminishes the efficiency when fewer paths would suffice. Consequently, designing a flexible graph index that supports arbitrary path combinations without index reconstruction remains a critical challenge.
@@
Challenge II: How to achieve efficient construction and effective search on the hybrid index? While Inner Product (IP) serves as the similarity metric (Bruch et al., 2024a; Wang et al., 2025), existing graph indexes optimized for it suffer from inefficient construction and redundant edge connectivity. Moreover, the indexing and querying processes of the hybrid index incur significant overhead and are inefficient on CPUs due to the massive hybrid distance calculations and the pruning of edges from various retrieval paths. Recent studies have demonstrated that accelerating index construction and query processing on GPUs is a promising solution (Ootomo et al., 2024; Li et al., 2025b). However, they do not apply to the hybrid index, where dense vectors and sparse vectors comparisons require vector dot product and set intersection operations, respectively. These disparate computational characteristics create challenges for efficient memory access and effective GPU parallelization. Furthermore, the computational demands of edges from various retrieval paths in the hybrid index introduce multiple efficiency bottlenecks. Semantic matching alone proves inadequate for complex queries. Although knowledge graphs provide complementary logical similarity (Zhou et al., 2026; Cao et al., 2025), a fundamental granularity mismatch exists: our index operates on document-level representations while knowledge graphs model fine-grained entity relationships. This disparity complicates the effective integration of logical information into the vector search process.
@@
To address these challenges, we propose Allan-Poe, a unified and flexible graph-based hybrid index accelerated by GPUs. Allan-Poe is built upon a unified lexical-semantic metric space, a framework that fuses diverse vector representations into a cohesive similarity metric, supporting retrieval under arbitrary fusion weights. To support this framework, we design an isolated heterogeneous edge storage mechanism that seamlessly integrates dense, sparse, and full-text retrieval, while optionally accommodating logic structures from knowledge graphs. To achieve efficient hybrid index construction, we develop a GPU-accelerated indexing pipeline featuring: (1) a warp-level hybrid distance computation kernel optimizing both dense and sparse operations parallelized on GPUs; (2) RNG-IP joint pruning that maintains search quality while reducing index complexity by combining Relative Neighborhood Graph (RNG) and Inner Product (IP) neighbor pruning; (3) keyword-aware neighbor recycling that preserves keyword search functionality by efficiently recycling the pruned neighbors to ensure keyword-based navigation on the index; and (4) an optional logical edge augmentation module that integrates the entity-level knowledge graph into the document-level hybrid index for complex queries. To deliver a high-performance search service, we design a dynamic query framework on GPUs that decouples computation from storage, supporting any combination of retrieval paths at runtime without requiring index reconstruction. The framework also incorporates: (1) dynamic heterogeneous edges loading for efficient traversal on the index; and (2) entity-document joint traversal for knowledge graph integration.
@@
  • Holistic hybrid indexing architecture. We propose a unified hybrid index that seamlessly integrates dense, sparse, and full-text retrieval paths with arbitrary weights, and optionally augment complex queries with knowledge graphs (§ 3).
  • Hardware-optimized construction pipeline. We present a novel GPU-accelerated indexing pipeline to overcome the inefficiency of hybrid computations and heterogeneous edge pruning with fine-grained parallelization (§ 4.1).
  • Dynamic and flexible query framework. We introduce a dynamic query framework that decouples computation from storage, enabling runtime path combinations with zero index reconstruction cost. (§ 4.2).
[…]
@@ 2.1. Related Work
(4) Knowledge graph search implements logical and semantic retrieval by converting queries into subgraphs and identifying similar structures through subgraph matching algorithms (Hu et al., 2017; Yuan et al., 2021; Sun et al., 2022). Subsequent enhancements incorporate entity and relation embeddings for improved efficiency (Zheng et al., 2020; Wang et al., 2017b). Recent approaches like GraphRAG leverage dense vector search and community detection for document retrieval of global questions (Edge et al., 2024; Cao et al., 2025; Zhou et al., 2026; Li et al., 2025a). While GraphRAG excels at summarization tasks, it typically underperforms vanilla RAG for simple question answering (Han et al., 2025). Unlike GraphRAG, our work selectively integrates logical information from knowledge graphs to enhance vector search, establishing a distinct paradigm applicable to more general scenarios beyond RAG.
@@
(1) Fusion retrieval integrates multiple retrieval methods within a unified index structure. Current fusion methods are primarily restricted to two-path combinations (Wang et al., 2025). For instance, DS-ANN (Zhang et al., 2024) employs a pre-defined fusion weight to combine dense and sparse vectors, constructing an HNSW index (Malkov and Yashunin, 2018) for efficient querying. While efficient, this method requires complete index reconstruction if the fusion weights change. IVF-Fusion (Bruch et al., 2024a) addresses this by reducing the dimensionality of sparse vectors before combining them with dense vectors, then using an IVF index for retrieval. Although this eliminates weight-dependent reconstruction, it limits the flexibility to select different retrieval paths for varying scenarios.
(2) Multi-path retrieval represents a more flexible paradigm that executes searches separately across different indexes and subsequently fuses the results (mil, 2025a; inf, 2025; ela, 2025). This paradigm is widely used in modern databases. However, the flexibility adversely affects query efficiency and accuracy while complicating index management.
@@ 2.2. Motivation
Recently, dense vector search has become the most popular paradigm in vector databases among single-path retrieval methods, distinguished by its capacity for comprehensive semantic representation (Pan et al., 2024; Wang et al., 2021a). Numerous specialized indexes have been developed to enhance their query efficiency and accuracy (Gao et al., 2025; Gou et al., 2025; Peng et al., 2023). However, vector similarity alone does not guarantee end-to-end relevance between queries and documents. To investigate this limitation empirically, we conduct experiments using two established real-world QA datasets: NaturalQuestions (NQ) (Wang et al., 2024b) for simple question answering and 2WikiMultiHopQA (WM) (Ho et al., 2020) for multi-hop question answering. For each dataset, we evaluate the first 1,000 queries, each associated with 1-3 ground-truth documents. We employ the BGE-M3 model (Chen et al., 2024) for embedding both documents and queries. The ground-truth baselines of vector similarity are established by computing the top-10 nearest neighbors via brute-force vector similarity search. We assess the retrieval accuracy of two state-of-the-art graph-based indexes: HNSW (CPU-based) (Malkov and Yashunin, 2018) and CAGRA (GPU-based) (Ootomo et al., 2024).
@@
As shown in Figure 2, while vector-similarity accuracy can easily reach 99% within 4ms on the CPU and 0.1ms on the GPU, the corresponding end-to-end accuracy is substantially lower. In practical document retrieval systems, it is this end-to-end accuracy—not vector-similarity accuracy—that dictates performance in downstream tasks. Furthermore, Figure 2 reveals that the end-to-end accuracy for the WM dataset is lower than that for NQ, whereas their maximal vector-based accuracies are comparable. This discrepancy underscores the limitations of dense vector search in handling complex queries. Consequently, reliance on single-path retrieval methods is insufficient for achieving satisfactory end-to-end performance, restricting their utility in downstream applications.
@@
Hybrid search has emerged as a powerful strategy to mitigate the limitations of single-path retrieval and improve end-to-end accuracy (inf, 2025; Wang et al., 2025). To evaluate its effectiveness, we measure the retrieval quality using Infinity (inf, 2025), a modern database featuring efficient hybrid search. To better assess the quality of retrieved documents, we employ Normalized Discounted Cumulative Gain at rank $k$ (nDCG@kk(Wang et al., 2013) with $k=5$, which evaluates both the recall and positional ranking of relevant documents in the retrieved results.
@@
Figure 3. Comparison of various retrieval paths on NQ using Infinity (inf, 2025). DVS, SVS, and FTS denote dense vector, sparse vector, and full-text search, respectively.
@@
However, no single configuration is optimal for all scenarios. Retrieval with more paths does not consistently outperform fewer or single paths. For example, in Figure 3(a), the two-path combination of dense vector and full-text search yields lower nDCG than the single-path methods using either dense or sparse vectors alone on dataset NQ. And the three-path combination shows lower nDCG than the two-path retrieval of dense and sparse vector search. On other datasets, the accuracy ranking may be completely different. Furthermore, different path combinations present a distinct trade-off between accuracy and efficiency. Although the single-path methods typically achieve lower accuracy than the multi-path approaches, they are more efficient and incur less overhead (Figure 3(b)), sometimes making it preferable for a few real-time applications. For instance, the sparse vector retrieval method achieves comparable accuracy with the three-path combination while maintaining lower latency. These findings underscore the necessity for flexible path combination in hybrid search. The trade-off strategies have been studied in (Wang et al., 2025), which is orthogonal to our work.
@@
Multi-path retrieval is widely adopted for flexible hybrid search in modern vector databases (inf, 2025; mil, 2025a; ela, 2025). Although it achieves superior accuracy compared to single-path approaches, this comes at the cost of increased time overhead. Moreover, this paradigm performs separate retrievals across individual indexes before fusing the results. The optimal results for a hybrid query may not be present within the top-$k$ results from any single path or may be excluded after fusing the results (Wang et al., 2024a, 2021b). To illustrate this, we examine the retrieval of top-5 documents from the NQ dataset using dense and sparse vectors independently. Figure 4 presents a representative example. Using dense vectors, ground-truth documents doc2 and doc4 are ranked 2nd and 30th, respectively, while sparse vectors rank them 3rd and 12th. If we fuse the top-5 results from both paths using equal weights (i.e., $0.5\timesdense$ $0.5\timessparse similarity$), doc4 will be excluded because it only involves the similarity score from the sparse path. To include doc4, more candidates should be included for each retrieval path (e.g., top-30). However, expanding candidates entails longer retrieval times, and it does not guarantee higher accuracy. This is because low-ranking noisy documents can be promoted during the fusion, disrupting the final accuracy.